A company is manufacturing building bricks and fire bricks. Both production require two processes. Brick forming and Heat treatment. The requirements for the two bricks are:

Building Bricks Fire Bricks
Forming per 100 bricks 3 hours 2 hours
Heating treatment per 100 bricks 2 hours 5 hours
Total cost of the two departments in one months were:
Forming $ 21 200
Heat treatment $ 48 800
Units produced during the month was: 130 000 70 000

Require:

Prepare statement of manufacturing costs for the two types of bricks.
Answer #1
1.1
The statement of manufacturing costs for the two types of bricks is prepared as below.
Statement Showing Manufacturing Costs of Building Bricks
Process Time for 100 Units (Hours) [A] Rate Per Hour ($) [B] Cost Per 100 Units [A*B] Cost for 130,000 Units
Brick Forming 3 4 $12 15,600 (130,000/100*12)
Heat Treatment 2 8 $16 20,800 (130,000/100*16)
$28 $36,400
Statement Showing Manufacturing Costs of Fire Bricks
Process Time for 100 Units (Hours) [A] Rate Per Hour ($) [B] Cost Per 100 Units [A*B] Cost for 70,000 Units
Brick Forming 2 4 $8 5,600 (70,000/100*8)
Heat Treatment 5 8 $40 28,000 (70,000/100*40)
$28 $33,600
Notes:
The rate per hour calculation is determined as follows:
Brick Forming:
Total Building Bricks Hours = (Building Bricks Produced during the Month*Number of Forming Hours Per 100 Bricks)/100 = (130,000*3)/100 = 3,900 hours
Total Fire Brick Hours = (Fire Bricks Produced during the Month*Number of Forming Hours Per 100 Bricks)/100 = (70,000*2)/100 = 1,400 hours
Total Hours Needed for Brick Forming = Total Building Brick Hours + Total Fire Brick Hours = 3,900 + 1,400 = 5,300 hours
Cost Per Hour of Brick Forming Department = Total Brick Forming Costs/Total Hours Needed for Brick Forming = 21,200/5,300 = $4 per hour
Heat Treatment:
Total Building Bricks Hours = (Building Bricks Produced during the Month*Number of Heat Treatment Hours Per 100 Bricks)/100 = (130,000*2)/100 = 2,600 hours
Total Fire Brick Hours = (Fire Bricks Produced during the Month*Number of Heat Treatment Hours Per 100 Bricks)/100 = (70,000*5)/100 = 3,500 hours
Total Hours Needed for Heat Treatment = Total Building Brick Hours + Total Fire Brick Hours = 2,600 + 3,500 = 6,100 hours
Cost Per Hour of Heat Treatment Department = Total Heat Treatment Costs/Total Hours Needed for Heat Treatment = 48,800/6,100 = $8 per hour
Answer #2
How to calculate rate per hour
answered by: Bittu Manna